OpenStudy (anonymous):

sin x (9-3)=2/5 What to do?

OpenStudy (skullpatrol):

Could you draw the equation please?

OpenStudy (skullpatrol):

$$sin x (9-3)=2/5$$

OpenStudy (anonymous):

Is it meant to be this: \[6\sin(x)=\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

\[\sin \alpha \left( 9-3 \right)=\frac{ 2 }{ 5 }\]

OpenStudy (anonymous):

Yes? you are right, but i can't do next step. Arcsin and so on

OpenStudy (anonymous):